Wild turkeys can see more clearly and in more detail than humans. Hence, turkey hunters need to dress themselves, and their weapons, in camouflage. Pretend that turkeys have a single gene, the A gene, which controls eyesight. (This is probably not really true, the pattern of inheritance is probably much more complicated.) Imagine that turkeys with a dominant allele (A) can see well while turkeys that are homozygous recessive cannot see as w wrell. A small population of 1000 wild turkeys has become isolated in a deep, Pennsylvania valley surrounded by mountains. These 1000 turkeys are physically and permanently separated from all of the other PA turkeys. Most of these 1000 turkeys cannot see well (Perhaps that is why they got separated from the other turkeys and cant get out of the valley.) The following table outlines the data from these turkeys. We will consider these turkeys the P generation. Complete any incomplete cells in the table. Genotypic Allelic Number (A) Allelic Number (a) Frequency Genotype Genotypic Number 2000 Aa 140 322 8001 .011 oO Total = Total = 1.0 Total A alleles Total a alleles- Total alleles = 2000 what is the allelie frequency ofA妨乍-au8y What is the allelic frequency of a(q) Using p & q.calculate the Expected genotypic frequencies for the next (F1) generation: AA = Aa = Ne Sot). (1322)-2A 29=2(·(oguD(314) 开A 門. 6f 506 20 니 - a.ã 个 2 2. aa = Calculate the Expected genotypic numbers for each genotype in the F1 generation of this population AA=
Answer #1

As per Hardy-Weinberg Equilibrium, let
p = Frequency of Dominant Allele
q = Frequency of Recessive Allele
Also p + q = 1
Now,
p2 = Frequency of Homozygous Dominant Genotype
2pq = Frequency of Heterozygous Dominant Genotype
q2 = Frequency of Homozygous Recessive Genotype

Here genotypic numbers of this population:
Number of Homozygous Dominant Genotype = p2 = 60
Number of Heterozygous Dominant Genotype = 2pq = 140
Number of Homozygous Recessive Genotype = q2 = 800

Genotypic frequency of Homozygous Dominant Genotype AA = Number / Total = 60/1000 = 0.06
​Genotypic frequency of Dominant Genotype Aa = Number / Total = 140/1000 = 0.14
​Genotypic frequency of Recessive Genotype aa = Number / Total = 800/1000 = 0.80

Allelic number (A) for Homozygous Dominant Genotype AA = AA Number*2 = 60*2 = 120
Allelic number (A) for Dominant Genotype Aa = Aa Number*1 = 140
Allelic number (A) for Recessive Genotype aa = aa Number*0 = 0 =
Total (A) = 260

Allelic number (a) for Homozygous Dominant Genotype AA = AA Number*0 = 0
Allelic number (a) for Dominant Genotype Aa = Aa Number*1 = 140
Allelic number (a) for Recessive Genotype aa = aa Number*2 = 800*2 = 1600
Total (a) = 1740

Allelic frequency of A = Total (A) / Total = 260/2000 = 0.13
Allelic frequency of a = Total (a) / Total = 1740/2000 = 0.87

Expected genotypic frequencies for next generation =
Genotypic frequency of AA = Allelic frequency of A * Allelic frequency of A = 0.13 * 0.13 = 0.0169
Genotypic frequency of Aa = 2 * Allelic frequency of A * Allelic frequency of a = 2 * 0.13 * 0.87 = 0.2262
Genotypic frequency of aa = Allelic frequency of a * Allelic frequency of a = 0.87 * 0.87 = 0.7569

Expected genotypic numbers for next generation (Assuming population of 10,000):
Genotypic Number of AA = Genotypic frequency of AA * 10000 = 0.0169 * 10000 = 169
Genotypic Number of Aa = Genotypic frequency of Aa * 10000 = 0.2262 * 10000 = 2262
Genotypic Number of aa = Genotypic frequency of aa * 10000 = 0.7569 * 10000 = 7569